# ChartMimic: Evaluating LMM’s Cross-Modal Reasoning Capability via Chart-to-Code Generation

### v1

@@ 4.4 Error Analysis
Figure 4: Error analysis of GPT-4V across four error types on the Direct Mimic task. The number in brackets indicates the count of error case. Error examples can be found in Appendix F.
We conduct a detailed error analysis of the state-of-the-art GPT-4V model to elucidate the current limitations of LMMs on ChartMimic, thereby identifying potential areas for further improvement. Our analysis identifies and category four main types of errors, detailed below.
Code-related Errors. (1) Dimension: Errors associated with data dimension, e.g., data dimension does not satisfy required conditions for operations; (2) Access: Errors occurring when accessing iterable elements out of bounds or accessing undeclared variables; (3) Parameter: Errors related to passing incorrect parameters when invoking functions. As observed in Fig. 4 (a), the majority of errors stem from Dimension, suggesting analysis and operation on data poses a challenge for GPT-4V.
Text-related Errors. (1) Detail: The text in the generated chart is largely consistent with the ground-truth chart text but contains minor discrepancies, such as mixing up “-” and “_”; (2) Missing: The generated chart omits text present in the ground-truth chart; (3) Extraneous: The generated chart contains text not found in the ground-truth chart, e.g., adding text for titles, reflecting the model’s own interpretation of the chart; As depicted in Fig. 4 (b), the majority of errors are Missing, which indicates that for GPT-4V, due to intensive information in charts, it is challenging for it to comprehend the whole scope of the charts, even the basic text recognition task.
Type-related Errors. (1) Confusion: GPT-4V misinterprets the chart type as one that appears or functions similarly, such as mistaking violin plots for box plots, and so on; (2) Missing: The generated chart omits chart types present in the ground-truth chart; (3) Extraneous: The generated chart includes chart types not present in the ground-truth chart. As demonstrated in Fig. 4 (c), although GPT-4V exhibits remarkable capability in object recognition for natural images, it still struggles with scientific charts, which contain more nuanced semantic meanings through visual logic.
Color-related Errors. (1) Similar: The colors are not the same as the ground-truth colors but appear analogous. (2) Different: The colors are entirely dissimilar to the ground-truth colors. As shown in Fig. 4 (d), though GPT-4V can not exactly recognize the accurate colors, it can identify similar ones.
To sum up, GPT-4V still faces challenges in code generation and exhibits notable visual understanding deficiencies on ChartMimic. It has difficulty accurately recognizing visual elements in figures or may hallucinate incorrect elements and struggle with complex data analysis. Combining these insights, there is a need for further improvement in both visual understanding and code generation.
@@ 5 Related Work
Large Multimodal Models. The proprietary LMMs such as GPT-4V [2], Gemini [3], and Claude-3 [28] have enabled complex multimodal interactions. Similarly, emerging open-weight LMMs such as LLaVA [33, 30], Qwen-VL [34], DeepSeek-VL [31] have contributed to the community. Despite these advancements, the effective evaluation of LMMs remains a major challenge. However, effectively evaluating LMMs remains a challenge, with open-source models performing well in benchmarks [12, 39, 40, 21, 41, 42, 43] yet falling short in practical applications [44, 45, 25]. This gap emphasizes the need for real-world-based evaluations. ChartMimic addresses this by requiring LMMs to translate complex visual information into code, testing their visual and coding capabilities.
@@
Code Generation. Tasks such as HumanEval [6], MBPP [7], APPS [8], and DS-1000 [9] are important benchmarks in natural language processing, but the inputs for these tasks are single-modal, consisting only of text. With the emergence of LMMs, the evaluation of multimodal code generation has become more important. On top of this, MMCode [24] attempts to address this issue, but their visual inputs may be simplistic and face the problem of a single evaluation method. Design2Code [25] is relevant to our work, and they evaluate LMMs’ code generation abilities through HTML web page generation. However, their test data comes from the C4 [27] dataset, which may pose a risk of data leakage. Meanwhile, they just imitate the HTML and do not take the customized instructions; we take this into consideration. Our ChartMimic provides a new set of manually curated $1000$ data pairs, ensuring chart diversity and offering more fine-grained evaluation methods.
@@
Chart Understanding. ChartMimic evaluates the capabilities of LMMs in grounding chart understanding into code generation. Previous works focus on chart question answering [21, 46, 39] and chart captioning [23, 22]. They assess the LMMs’ ability to understand specific data characteristics or summarize key information into text. ChartMimic advances the field by introducing a chart-to-code task, transforming the LMMs’ understanding of charts into code, which is neglected before but a realistic scenario for usage. This approach enables a comprehensive evaluation of the LMMs’ overall comprehension of charts and their ability to express this understanding in code form. Leveraging the linguistic properties of code, our benchmark introduces fine-grained metrics to assess LMMs’ chart understanding capabilities across multiple dimensions, including text, chart type, layout and color.
@@ 6 Conclusions
Appendix A Data Annotation
A.1 Data Annotation Principles
@@
Diversity of Chart Types. Data visualization has become an essential tool for conveying information in various fields, and the design practices and requirements for different types of charts vary significantly[47]. Most of previous work [48, 39, 21, 22, 23] focus only on line, bar, pie charts, etc, which are commonly used within the field of computer science. However, with the increasing integration of LMMs into everyday tasks, people from all fields are starting to use generative techniques as daily assistants to enhance their design and creative processes when creating visualizations. In this light, enriching the spectrum of chart diversity is crucial for evaluating LMMs’ proficiency in multimodal chart-to-code generation.
@@
Balance of Charts Complexity. Charts serve as visual aids for data presentation, enabling users to promptly grasp the underlying patterns and significance within the data. It is essential to adopt the appropriate chart type and complexity level to effectively convey the information [49]. Previous works have focused mainly on charts with only a single data format and low information density, which are rarely encountered in practical settings such as academic writing. Noting that LMMs like GPT-4V have already demonstrated outstanding data visualization capabilities, our focus is on using charts that are actually employed in practice, such as those from research papers, and on selecting charts with varying levels of complexity when constructing our benchmark.
@@
Reduction of Data Leakage. Recognizing that augmenting training data is a primary method for enhancing the performance of LMMs has led researchers to more comprehensively exploit all accessible data during pre-training. However, this approach introduces the possibility of data leakage, especially when pre-training data might already include resources such as the matplotlib gallery [^2] https://matplotlib.org/ or other pre-existing datasets, potentially resulting in inaccurate evaluations. To mitigate this issue, we deliberately avoid the use of code that can be readily found online or code that could be auto-generated by large language models for chart creation in constructing our dataset, thereby reducing the probability of data leakage. To further illustrate the potential data leakage in matplotlib gallery, we select the code for $20$ charts from the matplotlib gallery. Specifically, we provide the first half of the code as a prefix for Llama3-8B [15] and let the model complete the remaining code. Then we calculate the edit distance between the generated complete code block and the ground truth. Similarly, we apply the same process to the code in Direct Mimic, calculating the edit distance between the generated results and the ground truth. We find that the edit distance for the code in the matplotlib gallery is $22.1$, while the edit distance for Direct Mimic’s code is $39.8$. This indicates that the code in Direct Mimic have a larger edit distance, which further reduces the risk of data leakage compared to the matplotlib gallery ones.
@@
Integration of Authentic User Requirements. Users from diverse domains such as finance, healthcare, education, and engineering demonstrate unique needs and preferences for data visualization. These requirements extend beyond mere diversity; they demand the incorporation of charts capable of articulating complex and multi-dimensional data, and conforming to domain-specific aesthetic preferences[50, 49]. By aligning our dataset construction process with these real-world demands, we enable a more relevant and precise evaluation of LMMs. Adopting this approach not only reflects actual user patterns but also steers the research community in the iterative improvement of LMMs. This focused development will also lead to an enhanced user experience and increased user satisfaction, as the models become more proficient in meeting the sophisticated and varied needs of users.
A.2 Data Annotation Pipeline
@@
General Filtering. To obtain a high-quality dataset aligned with real-world use cases and to avoid data leakage, we initially scrape figures from source files of publications on arXiv [^3] https://arxiv.org/ that hold a CC BY 4.0 license and have a publication date after February 2024. We then extract figures in PDF format that contain at least one chart, yielding approximately $174,100$ figures across various domains (such as Physics, Mathematics, Computer Science, Quantitative Biology, Quantitative Finance, Statistics, Electrical Engineering and Systems Science, Economics). We further filter the figures based on criteria of how designers select inspiring visualization examples, including clarity and visual appeals[13, 14], color schemes[13], and uniqueness of the chart within its category, resulting in a refined collection of about $15,800$ charts. It is noteworthy that during this process, some charts, which do not entirely meet the clarity and aesthetic criteria, are retained in the corpus for their unique and rare qualities, pending further modifications.
Diversity and Information Density Filtering. In this phase, we adopt a two-stage filtering process focused on the diversity of the charts and the effectiveness of information communication, primarily assessed based on the complexity of visual elements and the information density of each chart[13]. Specifically, during the first round, we maintain a data pool, categorizing chart types and subcategories. Each chart is compared against the most similar one within the corresponding category in the data pool. If significant differences in visual complexity or information density are observed, the chart is added to the pool. Otherwise, we retain the one that better fulfills the criteria for diversity and effective information communication.
After the first round of filtration, we observe a decreasing frequency of charts with increasing levels of difficulty among the remaining $15,800$ entries. Charts with simpler designs and lower information density often share greater similarity in their arrangement of visual elements, leading to a higher rate of duplicates being identified and removed. In contrast, more complex charts with higher information density are less likely to have comparable counterparts, thereby reducing duplication and enhancing the retention rate of these complex charts. This process effectively refines the complexity spectrum of our dataset, ensuring a balanced distribution.
Subsequently, we conduct the second round of filtration. We engage five experts from different fields—computer science, visual communication design, and data visualization—to independently assess the data, creating five separate data pools. We then preserve the intersection of these pools and conduct a final, collective filtration from the combined set. The selection of which charts are to be included in the final pool is made through a voting process involving all five experts. Ultimately, this approach results in a curated collection of 279 charts.
@@
Manual Selection and Filling. In addition sourcing from arXiv, we curate chart figures from diverse platforms such as the matplotlib gallery, Stack Overflow, and plotting-related forums on Twitter and Reddit. These charts are deliberately chosen for their distinctive styles, which are not present in our arXiv dataset. To prevent data leakage, we meticulously process the data and styling of these charts, substituting existing style schemes and data with those not present in our dataset, while preserving their visual appeal.
@@
Code and Instruction Writing. We propose to manually write codes and instructions for ChartMimic based on the collected $500$ prototype charts. To ensure the quality of the annotation results, we assemble a team of skilled Python users. These individuals are master students from computer science-related disciplines, with at least 4 years of experience using Python and extensive familiarity with matplotlib. Initially, the annotators are tasked with meticulously reproducing the $500$ prototype charts using Python and the matplotlib package, resulting in a set of $500$ (figure, code, instruction) triplets for the Direct Mimic task. Subsequently, to simulate the real-world scenario of Customized Mimic, annotators are instructed to modify the chart data in the Direct Mimic task by integrating new data from various domains. They are then required to modify the corresponding codes and instructions, leading to the $500$ (figure, code, instruction) triplets for the Customized Mimic task. Consequently, we obtain the ChartMimic benchmark, comprising $1,000$ high quality test examples.
Appendix B Chart Taxonomy

### v2

@@ 4.4 Error Analysis
Figure 4: Error analysis of GPT-4o across four error types on the Direct Mimic task. The number in brackets indicates the count of error cases. Error examples can be found in Appendix Appendix I.
We conduct a detailed error analysis of the state-of-the-art GPT-4o model to elucidate the current limitations of LMMs on ChartMimic, thereby identifying potential areas for further improvement. Our analysis identifies and category four main types of errors, detailed below.
Code-related Errors. (1) Dimension: Errors associated with data dimension, e.g., data dimension does not satisfy required conditions for operations; (2) Access: Errors occurring when accessing iterable elements out of bounds or accessing undeclared variables; (3) Parameter: Errors related to passing incorrect parameters when invoking functions. As observed in Fig. 4 (a), the majority of errors stem from Dimension, suggesting analysis and operation on data poses a challenge for GPT-4o.
Text-related Errors. (1) Detail: The text in the generated chart is largely consistent with the ground-truth chart text but contains minor discrepancies, such as mixing up “-” and “_”; (2) Missing: The generated chart omits text present in the ground-truth chart; (3) Extraneous: The generated chart contains text not found in the ground-truth chart, e.g., adding text for titles, reflecting the model’s own interpretation of the chart; As depicted in Fig. 4 (b), the majority of errors are Missing and Extraneous, which indicates that for GPT-4o, due to intensive information in charts, it is challenging for it to comprehend the whole scope of the charts, even the basic text recognition task.
Type-related Errors. (1) Confusion: GPT-4o misinterprets the chart type as one that appears or functions similarly, such as mistaking violin plots for box plots, and so on; (2) Missing: The generated chart omits chart types present in the ground-truth chart; (3) Extraneous: The generated chart includes chart types not present in the ground-truth chart. As demonstrated in Fig. 4 (c), although GPT-4o exhibits remarkable capability in object recognition for natural images, it still struggles with scientific charts, which contain more nuanced semantic meanings through visual logic.
Color-related Errors. (1) Similar: The colors are not the same as the ground-truth colors but appear analogous. (2) Different: The colors are entirely dissimilar to the ground-truth colors. As shown in Fig. 4 (d), though GPT-4o can not exactly recognize the accurate colors, it can identify similar ones.
To sum up, GPT-4o still faces challenges in code generation and exhibits notable visual understanding deficiencies on ChartMimic. It has difficulty accurately recognizing visual elements in figures or may hallucinate incorrect elements and struggle with complex data analysis. Combining these insights, there is a need for further improvement in both visual understanding and code generation.
@@ 5 Related Work
Large Multimodal Models. The proprietary LMMs such as GPT-4o (OpenAI, 2024), Gemini (Google, 2023), and Claude-3 (Anthropic, 2024) have enabled complex multimodal interactions. Similarly, emerging open-weight LMMs such as LLaVA (Xu et al., 2024; Li et al., 2024a), InternVL (Chen et al., 2023), Qwen-VL (Bai et al., 2023), DeepSeek-VL (Lu et al., 2024) have contributed to the community. Despite these advancements, the effective evaluation of LMMs remains a major challenge. However, effectively evaluating LMMs remains a challenge, with open-source models performing well in benchmarks (Liu et al., 2023; Chen et al., 2024; Yue et al., 2023; Lu et al., 2023; Luo et al., 2025) yet falling short in practical applications (Xie et al., 2024; Koh et al., 2024; Si et al., 2024; Ji et al., 2024). This gap emphasizes the need for real-world-based evaluations that reflect authentic use cases (Ma et al., 2024a; b). ChartMimic addresses this by requiring LMMs to translate complex visual information into code, testing their visual and coding capabilities.
@@
Code Generation. Tasks such as HumanEval (Chen et al., 2021), MBPP (Austin et al., 2021), APPS (Hendrycks et al., 2021), and DS-1000 (Lai et al., 2023) are important benchmarks in natural language processing, but the inputs for these tasks are single-modal, consisting only of text, limiting their scope. With the emergence of LMMs, the evaluation of multimodal code generation has become increasingly critical for assessing real-world capabilities (Li et al., 2024b; Si et al., 2024; Wu et al., 2024a; Ge et al., 2025). For example, MMCode (Li et al., 2024b) attempts to address this issue, but their visual inputs may be simplistic and face the problem of a single evaluation method. Design2Code (Si et al., 2024) evaluates LMMs’ code generation abilities through HTML web page generation. However, their test data comes from the C4 (Raffel et al., 2020) dataset, which may pose a risk of data leakage. Meanwhile, they just imitate the HTML and do not take the customized instructions; we take this into consideration. Recently, Plot2Code (Wu et al., 2024a) undertakes work similar to ours, aiming to measure models’ code generation abilities through chart-to-code generation. Similarly, their approach of directly scraping data from the matplotlib gallery poses a risk of data leakage. Our ChartMimic provides a new set of manually curated $4,800$ data pairs, ensuring chart diversity and offering more fine-grained evaluation methods.
@@
Chart Understanding. ChartMimic evaluates the capabilities of LMMs in grounding chart understanding into code generation, bridging visual and programmatic domains. Previous works focus on chart question answering (Masry et al., 2022; Methani et al., 2020; Xu et al., 2023; Wang et al., 2024b; Li et al., 2024c; Zeng et al., 2024) and chart captioning (Rahman et al., 2023; Kantharaj et al., 2022). They assess the LMMs’ ability to understand specific data characteristics or summarize key information into text. ChartMimic advances the field by introducing a chart-to-code task, transforming the LMMs’ understanding of charts into code, which is neglected before but a realistic scenario for practical, real-world usage. This approach enables a comprehensive evaluation of the LMMs’ overall comprehension of charts and their ability to express this understanding in code form. Leveraging the linguistic properties of code, our benchmark introduces fine-grained metrics to assess LMMs’ chart understanding capabilities across multiple dimensions, including text, chart type, layout and color.
@@ 6 Conclusions
Acknowledgments
This work was partly supported by the National Key Research and Development Program of China (No. 2024YFB2808903) and the Shenzhen Science and Technology Program JSGG20220831110203007).
@@ Acknowledgments
Appendix A Author Contributions
Yujiu Yang served as the principal advisor for this project. Cheng Yang and Chufan Shi acted as the student project leads.
Data Annotation. Yaxin Liu and Bo Shui co-led the data annotation process (Appendix B). Chufan Shi, Cheng Yang, Mohan Jing and Linran Xu helped to contribute to various aspects of this process. Xiaomei Nie provided expert guidance on data visualization and conducted quality control.
Code Implementation. Cheng Yang and Chufan Shi co-implemented the codebase for ChartMimic.
Experiments. Cheng Yang and Chufan Shi jointly conducted the evaluation of LMMs on ChartMimic. Yaxin Liu and Bo Shui jointly led the human evaluation.
Paper Writing. Cheng Yang and Chufan Shi finished the main manuscript. Junjie Wang and Yujiu Yang provided in-depth reviews and revisions. Yaxin Liu and Bo Shui created all visualizations presented in the paper and contributed to the Data Annotation (Appendix B) and Chart Taxonomy (Appendix D). Deng Cai, Xinyu Zhu, Siheng Li, Yuxiang Zhang and Gongye Liu contributed through proofreading, discussions and feedback.
Appendix B Data Annotation
B.1 Data Annotation Principles
@@
Diversity of Chart Types. Data visualization has become an essential tool for conveying information in various fields, and the design practices and requirements for different types of charts vary significantly(Parsons, 2021). Most of previous work (Han et al., 2023; Xu et al., 2023; Masry et al., 2022; Kantharaj et al., 2022; Rahman et al., 2023) focus only on line, bar, pie charts, etc, which are commonly used within the field of computer science. However, with the increasing integration of LMMs into everyday tasks, people from all fields are starting to use generative techniques as daily assistants to enhance their design and creative processes when creating visualizations. In this light, enriching the spectrum of chart diversity is crucial for evaluating LMMs’ proficiency in multimodal chart-to-code generation.
@@
Balance of Charts Complexity. Charts serve as visual aids for data presentation, enabling users to promptly grasp the underlying patterns and significance within the data. It is essential to adopt the appropriate chart type and complexity level to effectively convey the information (Evergreen, 2019). Previous works have focused mainly on charts with only a single data format and low information density, which are rarely encountered in practical settings such as academic writing. Noting that LMMs like GPT-4o have already demonstrated outstanding data visualization capabilities, our focus is on using charts that are actually employed in practice, such as those from research papers, and on selecting charts with varying levels of complexity when constructing our benchmark.
@@
Reduction of Data Leakage. Recognizing that augmenting training data is a primary method for enhancing the performance of LMMs has led researchers to more comprehensively exploit all accessible data during pre-training. However, this approach introduces the possibility of data leakage, especially when pre-training data might already include resources such as the matplotlib gallery [^2] https://matplotlib.org/ or other pre-existing datasets, potentially resulting in inaccurate evaluations. To mitigate this issue, we deliberately avoid the use of code that can be readily found online or code that could be auto-generated by large language models for chart creation in constructing our dataset, thereby reducing the probability of data leakage. To further illustrate the potential data leakage in matplotlib gallery, we select the code for $20$ charts from the matplotlib gallery. Specifically, we provide the first half of the code as a prefix for Llama3-8B (AI@Meta, 2024) and let the model complete the remaining code. Then we calculate the edit distance between the generated complete code block and the ground truth. Similarly, we apply the same process to the code in Direct Mimic, calculating the edit distance between the generated results and the ground truth. We find that the edit distance for the code in the matplotlib gallery is $22.1$, while the edit distance for Direct Mimic’s code is $39.8$. This indicates that the code in Direct Mimic have a larger edit distance, which further reduces the risk of data leakage compared to the matplotlib gallery ones.
@@
Integration of Authentic User Requirements. Users from diverse domains such as finance, healthcare, education, and engineering demonstrate unique needs and preferences for data visualization. These requirements extend beyond mere diversity; they demand the incorporation of charts capable of articulating complex and multi-dimensional data, and conforming to domain-specific aesthetic preferences(Qin et al., 2020; Evergreen, 2019). By aligning our dataset construction process with these real-world demands, we enable a more relevant and precise evaluation of LMMs. Adopting this approach not only reflects actual user patterns but also steers the research community in the iterative improvement of LMMs. This focused development will also lead to an enhanced user experience and increased user satisfaction, as the models become more proficient in meeting the sophisticated and varied needs of users.
B.2 Data Annotation Pipeline
Figure 5: An illustration of the data annotation pipeline, which encompasses five key steps.
We present the detailed description about the five-step data annotation pipeline in this section. Fig. 5 demonstrates an illustration of the data annotation pipeline.
@@ B.2 Data Annotation Pipeline
General Filtering. To obtain a high-quality dataset aligned with real-world use cases and to avoid data leakage, we initially scrape figures from source files of publications on arXiv [^3] https://arxiv.org/ that hold a CC BY 4.0 license and have a publication date after February 2024. We then extract figures in PDF format, yielding approximately 174,100 figures across various domains (such as Physics, Mathematics, Computer Science, Quantitative Biology, Quantitative Finance, Statistics, Electrical Engineering and Systems Science, Economics). We filter these figures based on file format and generation method, retaining only Matplotlib-generated PDFs, indicating that these figures can be reproduced using Python. This process results in a refined collection of 15,800 figures.
Diversity and Information Density Filtering. This stage involves a two-phase process conducted by five domain experts from information visualization, digital media, industrial design, visual communication design, and computer science. The information visualization expert has three years of research experience in scientific visualization and visual analytics. The digital media specialist focuses on human-computer interaction and multimedia design. The industrial design expert brings perspectives from user experience and product visualization. The visual communication design expert specializes in graphic design principles and information aesthetics. The computer science expert has extensive experience in data visualization programming and scientific computing.
In the first phase, these experts conduct a 7-day manual review of the 15,800 figures, focusing on visual diversity and information communication effectiveness. They reference the Matplotlib gallery, gradually identifying and finalizing chart type while reviewing the figures, and build corresponding type pools. For each new figure, they assess its visual elements—such as layout, axes, line styles, marker styles, and colors—against existing figures in their corresponding type pool. If the figure as long as exhibits a distinctive difference in at least one of these aspects, it is retained; otherwise, it is excluded. This process results in 1,295 figures being selected for the second phase.
In the second phase, these experts independently review the 1,295 figures and further select those figures they deem to exhibit significant distinctions and diversity. Figures selected unanimously by all experts are directly included, while the remaining figures are subjected to a majority voting system requiring at least 3/5 votes for inclusion. This rigorous process, which takes less than 3 days to complete, results in a final set of 279 figures.
@@
Manual Selection and Filling. In addition to sourcing from arXiv, we curate chart figures from diverse platforms such as the matplotlib gallery, Stack Overflow, and plotting-related forums on Twitter and Reddit. These charts are deliberately chosen for their distinctive styles, which are not present in our arXiv dataset. Consequently, we obtain 600 prototype charts for ChartMimic. This stage took us less than a week to complete the data selection.
@@
Code and Instruction Writing. We propose to manually write codes and instructions for ChartMimic based on the collected 600 prototype charts. To ensure annotation quality, a team of skilled Python users—Python annotators—master’s students in computer science with 6+ years of Python and matplotlib experience—reproduce 600 prototype charts using Python 3.9.0 and matplotlib v3.8.4. Since the unannotated data in the figures cannot be fully restored, they can only be approximated when writing the code. This process generates 600 (figure, code, instruction) triplets for the Direct Mimic task and another 600 triplets for the Customized Mimic task by integrating data from various domains into the corresponding code and instructions, comprising 1,200 high-quality seed data.
Data Augmentation. After developing the seed triplets, we proceed with a manual data augmentation process. Python Annotators modify various elements of each seed triplet, including data, color schemes, and mark styles, to create augmented triplets. For each seed triplet, we generate three additional augmented triplets. This process enhances our dataset, resulting in a total of $4,800$ triplets. The "Code and Instruction Writing" and "Data Augmentation" stages together take the data annotators approximately 1.5 months to complete.
B.3 Complexity Levels
Our complexity assessment framework is established upon two fundamental criteria, designed to comprehensively evaluate both the visual and technical aspects of chart generation:
Visual Elements Complexity We systematically evaluate the sophistication of each chart through a comprehensive scoring mechanism applied to its constituent visual components. This encompasses the chart typology, data grouping structures, marker configurations, textual elements, chromatic schemes, compositional layouts, and coordinate systems. Each visual component is quantitatively assessed based on both its frequency and sophistication level (designated as 1-3 points for low/medium/high complexity, respectively). This multi-dimensional scoring approach ensures a thorough evaluation of the visual complexity inherent in each chart.
Implementation Complexity We incorporate code complexity as a quantitative metric, measured primarily through code length and structural intricacy. This parameter effectively captures the technical sophistication required for accurate chart reproduction, including the complexity of data preprocessing, visualization logic, and stylistic customizations. The implementation complexity provides insights into the programming challenges associated with each chart type.
Consequently, charts in ChartMimic are systematically categorized into three distinct complexity levels, each representing a specific combination of visual and implementation challenges:
Easy: Fundamental chart configurations featuring minimal visual complexity and straightforward implementation requirements (e.g., in LABEL:fig:easy-case, monochromatic bar charts with sparse textual elements and simplified data representation).
Medium: Charts exhibiting intermediate complexity in visual element composition or implementation requirements (e.g., in LABEL:fig:medium-case, dual-subplot bar charts incorporating grouped data structures, diverse chromatic schemes, and moderate textual annotations).
Hard: Charts demonstrating sophisticated visual elements or advanced implementation (e.g., in LABEL:fig:hard-case, complex dual-subplot bar charts featuring divergent data patterns, extensive color schemes, comprehensive textual annotations, and substantial code complexity).
B.4 Instruction Examples
To illustrate our instruction format and task requirements, we present examples (bar_28 and CB_29) with their corresponding (figure, instruction, code) triplets for both Direct Mimic and Customized Mimic tasks. Fig. 8 and Fig. 9 show the Direct Mimic and Customized Mimic tasks for a bar chart (bar_28), respectively. The instruction provides guidance on creating the visualization. Similarly, Fig. 10 and Fig. 11 demonstrate the tasks for a more complex combination chart (CB_29).
Figure 8: An Example of the Direct Mimic task (bar_28), showing the (figure, instruction, code) triplet. Additionally, we also display the ground truth figure rendered from the code for illustration.
Figure 9: An Example of the Customized Mimic task (bar_28), showing the (figure, instruction, code) triplet. Additionally, we also display the ground truth figure rendered from the code for illustration.
Figure 10: An Example of the Direct Mimic task (CB_29), showing the (figure, instruction, code) triplet. Additionally, we also display the ground truth figure rendered from the code for illustration.
Figure 11: An Example of the Customized Mimic task (CB_29), showing the (figure, instruction, code) triplet. Additionally, we also display the ground truth figure rendered from the code for illustration.
Appendix C Correlation Between Test Set and Testmini set
Table 7: The ChartMimic benchmark with Direct Mimic task. We report the results of both Testmini and Test set.
Model Params Test set Exec. Rate Low-Level High-Level Overall
Text Layout Type Color Avg. GPT-4o
Proprietary
GPT-4o - Testmini 93.2 81.5 89.8 77.3 67.2 79.0 83.5 81.2
Test 93.0 83.4 90.2 76.7 66.0 79.1 84.1 81.6
Open-Weight
InternVL2-2B 2.2B Testmini 52.5 23.6 35.8 16.0 15.4 22.7 24.2 23.5
Test 51.8 23.3 34.8 15.9 15.9 22.5 25.6 24.0
InternVL2-8B 8.1B Testmini 61.8 31.5 51.1 28.6 26.2 34.4 38.9 36.6
Test 61.6 33.4 47.4 25.7 24.2 32.7 39.9 36.3
InternVL2-26B 26.0B Testmini 69.3 39.2 58.7 35.9 31.8 41.4 47.4 44.4
Test 69.9 41.7 58.0 35.6 31.0 41.6 48.1 44.9
InternVL2-Llama3-76B 76.0B Testmini 83.2 54.1 74.5 49.2 41.5 54.8 62.2 58.5
Test 83.3 55.6 73.5 50.4 40.8 55.1 62.7 58.9
Table 8: The ChartMimic benchmark with Customized Mimic task. We report the results of both Testmini and Test set.
Model Params Test set Exec. Rate Low-Level High-Level Overall
Text Layout Type Color Avg. GPT-4o
Proprietary
GPT-4o - Testmini 96.5 88.5 92.9 79.2 67.6 82.1 84.3 83.2
Test 96.2 87.2 91.7 80.1 66.4 81.4 84.8 83.1
Open-Weight
InternVL2-2B 2.2B Testmini 49.3 22.2 35.4 20.0 18.1 23.9 27.8 25.9
Test 49.6 22.4 33.9 19.2 19.6 23.8 28.4 26.1
InternVL2-8B 8.1B Testmini 73.0 43.1 54.4 39.9 35.4 43.2 48.9 46.1
Test 73.5 43.7 54.1 41.1 34.1 43.3 49.8 46.5
InternVL2-26B 26.0B Testmini 73.7 43.9 62.3 43.5 34.3 46.0 51.1 48.6
Test 74.7 44.7 66.3 46.8 35.1 48.2 50.8 49.5
InternVL2-Llama3-76B 76.0B Testmini 89.8 57.8 79.0 63.5 50.5 62.7 66.7 64.7
Test 88.1 57.9 79.6 65.6 51.7 63.7 68.2 66.0
Tab. 7 and 8 reports the performance of GPT-4o and InternVL2 series LMMs (2B, 8B, 26B, Llama3-76B) on Direct Mimic and Customized Mimic task. The minor differences between scores on the test subset and the testmini subset suggest that testmini effectively mirrors the test subset, serving as a valuable evaluation subset for model development, especially for those who have limited computing resources.
Appendix D Chart Taxonomy